OpenStudy (anonymous):

Sasha solved an equation. Her work is shown below. Explain Sasha's error. Then, show the correct way to solve. h divided by seven equals six. Add seven to both sides: h divided by seven plus seven equals six plus seven, so h equals thirteen.

OpenStudy (anonymous):

h/7 = 6 (h/7)+7 = 13 the wrong is to add 7 no need for that you need to multiply by 7 h/7=6 h=7*6=42

OpenStudy (anonymous):

the inverse operation of division is multiplication,while addition is the inverse of subtraction.